# Video-based Automatic Lameness Detection of Dairy Cows using Pose Estimation and Multiple Locomotion Traits

### v1

@@ 1 Introduction
Lameness is a painful gait disorder in dairy cows and is often characterized by abnormal locomotion of the cow. A recent literature review (thomsen2023prevalence, ) estimated the global prevalence of lameness at 22.8%, with little change in the last 30 years. Lameness has a negative impact on welfare (whay2017impact, ) and leads to substantial economic losses (enting1997economic, ) due to decreased milk production and reproduction (huxley2013impact, ) as well as premature culling (enting1997economic, ). While lameness is commonly assessed by trained observers performing visual locomotion scoring of the herd, the procedure is time-consuming and cannot realistically be performed on a regular basis. Hence, dairy farms could benefit from automatic lameness detection.
@@
To date, a number of studies have investigated ways to automate locomotion scoring and lameness detection using camera systems. Video cameras are an attractive sensor for this application as they are relatively inexpensive, non-intrusive, and scale well with large herds. A three-step approach is commonly taken to detect lameness from videos: (1) use computer vision methods to localize body parts of interest, (2) compute one or more locomotion traits from the extracted body parts, and (3) train a classifier to score lameness using the locomotion traits as features. In the past, the body parts were localized using classical computer vision methods such as background subtraction (song_automatic_2008, ; poursaberi_real-time_2010, ; zheng_cows_2023, ; zhao_automatic_2023, ). These methods worked in experimental settings but were sensitive to changes in background and light, making them less applicable in practice. Others placed physical markers (tags or paint marks) on the cows’ body parts and tracked the markers with specialized software blackie_associations_2013 ; karoui_deep_2021 . In practical settings, however, physical markers don’t scale well to large herds as they need to be placed on each cow and cleaned regularly to remain visible. More recently, with the emergence of deep neural networks, studies started using deep-learning-based object detection wu_lameness_2020 ; kang_accurate_2020 ; jiang_dairy_2022 ; zheng_cows_2023 to localize the legs or the back of the cows, object segmentation arazo_segmentation_2022 to extract the body contour from the background, or markerless (i.e., without physical markers) pose estimation mathis_deeplabcut_2018 ; russello_t-leap_2021 ; zhao_automatic_2023 ; barney2023deep ; taghavi2023cow to localize multiple body parts in videos. Although they typically require more data than classical approaches, the deep-learning methods cope well with complex background and light conditions and can sometimes even cope with occlusions such as fences russello_t-leap_2021 ; taghavi2023cow .
@@
Once localized in the images or video frames, the outline of the spine, for instance, can be used to compute the back posture poursaberi_real-time_2010 ; viazzi_analysis_2012 ; van_hertem_automatic_2014 ; viazzi_comparison_2014 ; van_hertem_implementation_2018 ; jiang_dairy_2022 ; zheng_cows_2023 , and the location of the legs to compute the tracking distance song_automatic_2008 ; blackie_associations_2013 or stride length blackie_associations_2013 ; wu_lameness_2020 ; zheng_cows_2023 . To the best of our knowledge, almost all studies on lameness detection from videos use only one locomotion trait as a feature to score lameness, and only two studies barney2023deep ; zhao_automatic_2023 combined two locomotion traits, namely back posture and head bobbing.
@@
Using the locomotion trait(s) as feature(s), supervised learning classifiers can then be trained to score lameness. In supervised learning, classifiers learn from given examples, also known as ground truth or golden standard. Manual locomotion scores, that is, locomotion scores provided by one or more observers, make up the ground truth of lameness detection classifiers. The subjective nature of manual locomotion scoring is a well-known problem (schlageter-tello_effect_2014, ) and often leads to low intra- and inter-observer reliability and agreement. However, a classifier can only be as good as its ground truth, so information about the reliability of the locomotion scale is necessary. However, observer reliability and agreement are seldom reported, let alone analyzed.
@@
Three critical gaps emerge from the studies discussed so far: (1) the use of obsolete image processing methods remains frequent, (2) no one combined more than two locomotion traits for lameness classification, and (3) the reliability of the ground truth is seldom reported. This paper addresses the three gaps mentioned above and proposes a non-intrusive and fully automated approach to camera-based lameness detection that includes multiple locomotion traits.
@@
We used videos of walking cows that were scored on a 5-point locomotion scoring scale by four observers. We first reported and discussed the intra- and inter-observer reliability and agreement of the ground truth. We showed the effect of several approaches for merging scores from multiple observers and motivated the merging of the 5-point locomotion scale to a binary scale. We then trained T-LEAP russello_t-leap_2021 , a deep-learning markerless pose estimation model, to automatically extract the motion of multiple body parts (later referred to as keypoints) from videos of walking cows. The sequences of keypoints were used to compute six locomotion traits that are known to be correlated with locomotion scores schlageter-tello_relation_2015 , namely back posture measurement, head bobbing, tracking distance, stride length, stance duration, and swing duration. Using the locomotion traits mentioned above as input features, we trained multiple machine-learning classifiers to score the gait on a 2-level scale (healthy/lame). We evaluated the performance of each model and showed the impact of using different combinations of locomotion traits on the score classification.
@@ 2.1 Data acquisition
The data were collected in Tilburg, The Netherlands, at a commercial dairy farm whose herd contained about a hundred Holstein-Frisian cows. The data were collected between 9 am and 4 pm on 8 different days between May and July 2019. The cows were filmed from the side while they walked freely through an outdoor passageway. A ZED RGB-D stereo camera [^1] https://www.stereolabs.com/zed-2/ was placed 2 meters above the ground, at 4.5m from the fence of the passageway. The camera directly faced the passageway and recorded in landscape mode at Full-HD (1080p) resolution at 30 frames per second. The recordings were saved into short videos of about 7.6 seconds, which was the average time a cow needed to walk the visible part of the passageway (9.5 meters). The same data acquisition campaign was used by russello_t-leap_2021 on the same farm. In total, 1101 videos were collected, and a subset of 272 videos were selected according to the following criteria: there was only one cow on the passageway, and the cow walked from the left to the right without distraction or interruption.
@@ 2.2 Locomotion scoring
The locomotion scoring was performed using the 5-point discrete scale described by Sprecher et al. 1997 sprecher_lameness_1997 , where a score of 1 corresponds to normal gait, 2 to midly lame, 3 to moderately lame, 4 to lame and 5 to severely lame. The videos were scored by four observers: one expert (A) with 20 years of experience in visual locomotion scoring and 3 observers (B, C, D) with no prior experience in locomotion scoring but with a background in animal science and dairy farming. The inexperienced observers were trained by the expert (A) before the scoring session. During the scoring session, each video was played twice in a row to give enough time to observe the locomotion. To ensure consistency, the observers were asked to give the lowest score if they were hesitating between two scores. All the videos were scored on the same day. After the scoring session, the observers indicated no cow recognition, i.e., that they did not recognize the individual cows that appeared in multiple videos. Table 1 shows the distribution of the scores assigned by the four observers. The distribution of the scores was highly imbalanced and indicated a homogeneous herd, where most cows were distributed throughout the first two levels of the scale (normal, mildly lame), which is typical of herds with a low prevalence of lameness thomsen2008evaluation .
@@ 2.3 Observers reliability and agreement
Manual locomotion scoring is subjective flower_effect_2006 . Investigating the reliability and agreement between (inter-rater) and among (intra-rater) raters can inform on the quality of the data. Reliability estimates the capability of the raters to differentiate between the different scores, whereas agreement assesses the capability of the raters to assign the same score to the same data point. Reliability was measured with Krippendorff’s $\alpha$ krippendorff2011computing for ordinal values, and agreement was presented as the Percentage of Agreement (PA) and Specific Agreement (SA). The commonly accepted thresholds are $\alpha\geq 0.66$ for reliability krippendorff2011computing , and $\text{PA}\geq 75\%$ for agreement schlageter-tello_effect_2014 . The inter-observer and intra-observer measures are reported in the following sub-sections.
2.3.1 Inter-observer reliability and agreement
The inter-observer reliability and agreement values are reported in Table 2. The $\alpha$ value was marginally lower than the commonly accepted threshold. It meant that the observers agreed on 60% of the labels they were expected to disagree on by chance. The percentage of agreement was also low. When looking at the specific agreement, score 5 had the lowest agreement. Observer A didn’t assign any score of 5, whereas the other observers assigned a score of 5 to at most three videos, hinting that the boundary between 4 and 5 was not clear to the inexperienced observers.
Table 2: Inter-observer reliability ($\alpha$), agreement (PA), and agreement per locomotion score (specific agreement) on the 5-point locomotion scale.
Levels $\alpha$ PA Specific Agreement
1 2 3 4 5
1-2-3-4-5 0.602 55.8 69.7 49.4 37.0 44.4 28.6
2.3.2 Intra-observer reliability and agreement
Intra-observer metrics are usually performed on repeated ratings from the same observer on the same data points. Here, however, the videos were only scored once, so we could not compute intra-observer metrics the usual way. Instead, we proposed the following approach to approximate the intra-observer reliability and agreement. As mentioned in sub-section 2.1, some cows were present in several videos of the dataset. Assuming that the locomotion score remained the same for a period of time $T$, we could consider videos of a cow recorded less than $T$ hours apart to be the same data sample and should, therefore, be assigned the same score by the observers. We set $T=48$ hours and found 55 pairs of videos of the same cows recorded at less than 48-hour intervals. This data was then used to approximate the intra-observer metrics. We would like to emphasize that here, the intra-observer metrics were approximated because they were computed on a subset of the scores.
The intra-observer reliability and agreement values are reported per observer in table 3. Out of the four observers, only observer A and observer C had the highest $\alpha$, meaning that these observers were the best at distinguishing between the different levels of the scale. None of the observers reached an acceptable level of agreement, meaning that they gave the same score to the same cow less than 75% of the time.
Table 3: Intra-observer reliability ($\alpha$), agreement (PA), and agreement per locomotion score (specific agreement) on the 5-point locomotion scale.
Observer $\alpha$ PA Specific Agreement
1 2 3 4 5
A 0.611 56.4 72.0 46.2 20.0 54.5 0.0
B 0.552 49.1 71.2 9.5 22.2 40.0 100.0
C 0.653 60.0 72.3 60.9 36.4 0.0 0.0
D 0.585 58.2 76.9 32.0 30.8 33.3 0.0
@@ 2.4 Merging the locomotion scores
The locomotion scores ranged from 1 to 5, and were provided by multiple observers. Our task at hand, however, was a binary classification task, where the model was taught to distinguish between normal and lame gaits based on ground-truth examples. The ground-truth consisted of one binary label (normal/lame) per sample (video). Therefore, the locomotion scores needed to be merged in two ways: first, the scores from the multiple observers needed to be merged into one value; second, the five levels of the scale needed to be merged into a binary scale.
2.4.1 Merging the scores from multiple observers
For a classification task, each sample (i.e., video) is assigned one ground-truth label or locomotion score based on the multiple ground-truth labels provided by the observers. Common strategies for merging scores from multiple observers are mean, majority voting, and weighted voting. In the case of a tie with voting, the highest or the lowest score is retained. A drawback of these merging strategies is that if one or more observers have low reliability and agreement, chances are that their contributions would still add noise to the ground truth. Using the scores of only one observer, e.g., the most reliable observer, could also be a valid strategy, but one is taking the risk of training the classifier with observer bias. We therefore proposed an additional merging strategy: $\tau$-voting, where $\tau$ defined a minimum reliability threshold. The scores of an observer were then included in the vote if its intra-observer reliability was $\geq\tau$. We set the threshold to the overall inter-observer reliability on the 5-level scale, so $\tau=0.602$.
Table 4 shows the intra-reliability and intra-agreement values after applying the different merging strategies. The $\tau$-vote strategy increased both metrics the most, where only scores provided by the two most reliable observers (A and C) were included in the majority voting. Because there were only two observers included, majority voting was here equivalent to taking the lowest value of the two scores upon disagreement. This approach aligned with the direction given in the scoring session to assign the lowest score if an observer is uncertain. As shown in table 4, merging the scores largely improved the agreement compared to the individual observers and brought the reliability above the acceptable threshold when using majority voting and $\tau$-voting. As a result, the locomotion scores were merged into one ground-truth value using the votes of observers A and C with $\tau$-voting.
Table 4: Intra-observer reliability and agreement of the different voting strategies used for merging the scores from multiple observers.
Voting strategy $\alpha$ PA Specific Agreement
1 2 3 4 5
Mean 0.614 58.2 68.1 53.3 22.2 66.7 0.0
Weighted vote 0.611 56.3 72.0 46.1 20.0 54.5 0.0
Majority vote 0.667 65.4 82.3 38.5 22.2 57.1 0.0
$\tau$-vote ($\tau=0.6$) 0.695 70.9 83.1 58.1 44.4 40.0 0.0
2.4.2 Merging the levels of the scale
@@
The majority of the studies on lameness detection focus on 2-level (normal, lame) or 3-level (normal, moderately lame, lame) locomotion scores rather than 5-level scale sprecher_lameness_1997 ; flower_effect_2006 . There are two primary motivations for resorting to smaller resolutions in locomotion scores. First, severely lame cows are rare to find, as most of them get treatment or are culled before they reach this level of lameness engel_assessment_2003 . This results in a heavily unbalanced score distribution, most scores being levels 1, 2, and 3. It is then challenging to train a classifier on unbalanced datasets, especially when little examples are available for some classes. Second, visual locomotion scoring is subjective and often yields low intra- and inter-observer agreement and reliability measures. schlageter-tello_effect_2014 studied the effects of merging the levels of the locomotion scoring scale and showed that while the agreement and reliability measures were shown to be low for 5-level scales, they only exceeded the acceptable threshold for 2-level scales. We then followed the same practice as schlageter-tello_effect_2014 , and merged our 5-level scale to a 2-level (i.e., binary) scale, where level 1 indicates a normal gait, and levels 2,3,4 and 5 indicate a lame gait. The levels of the scale were merged into a binary scale after merging the scores from the multiple observers. This resulted in an intra-observer agreement of 80%, and reliability of 0.590. Note that reliability metrics such as Krippendorff’s $\alpha$ can decrease when the scoring scale is smaller because the chance of agreement is larger.
\@fb@secFB
2.5 Overview of the Materials
To summarize, the data used for this study consisted of 272 videos of walking cows, with 98 unique cows. For each video, there was one binary ground-truth label or locomotion score. In total, 143 videos were labeled as normal, and 129 videos were labeled as lame.
@@ 3 Methods
Our methodology consisted of three main parts: pose estimation, gait features extraction, and gait classification. These parts are described in detail in the following subsections, and a graphical summary of the methods is provided in Figure 1.
@@ 3.1 Pose estimation
Pose estimation models can be used to predict the position of keypoints (body parts) in images and videos without requiring physical markers. T-LEAP is a recent, deep-learning-based, temporal pose estimation model that was trained to detect keypoints on the body of cows in videos russello_t-leap_2021 . The model used sequences of successive frames to predict the coordinate of the keypoints, and was shown to perform better than static approaches in the presence of occlusions (such as fences). In this study, we used T-LEAP to extract nine keypoint coordinates from the video frames (Figure 2). In the next paragraphs, we describe the steps necessary for image cropping, pose estimation, and correction.
@@ 3.1.1 Detect-and-crop
The T-LEAP model required the input frames to be square and cropped around the cow’s body. The cows were automatically localized in the video frames using the Faster Region-based Convolutional Neural Network (Faster R-CNN), an object-detection model that returns the coordinates of a bounding box (bbox) around each object of interest (here, cows). We used the Faster R-CNN model (with ResNeXt-101 backbone) trained on the COCO-2017 dataset from the Detectron2 library wu2019detectron2 . The COCO-2017 dataset contained 118K training images with annotations for 80 categories of objects, among which 8014 bounding-box annotations of cows. The Faster R-CNN model from Detectron2 worked out of the box and could detect the cows in our video frames without fine-tuning. Each frame of each video was fed to the object-detection model, which returned a list of bounding boxes, one for each detected cow. For each frame, the bounding box was made square by extending the top and bottom coordinates to match the width while keeping the cow vertically centered. A 100-pixel padding was added to all four sides to ensure that the body of the cow was fully visible in the cropped area. The image was cropped to the coordinates of the extended bounding box and re-scaled to a size of $200\times 200$ pixels. The coordinates of the cropping bounding box were saved to transform the keypoint predictions back to the true coordinates for the video frame.

### v2

@@ Abstract
keywords:
lameness, detection, cows, locomotion, pose-estimation, deep-learning
@@ 1 Introduction
Lameness is a painful gait disorder in dairy cows and is often characterized by abnormal locomotion of the cow.
@@
A recent literature review [1] estimated the global prevalence of lameness at 22.8%, with little change in the last 30 years. Lameness has a negative impact on welfare [2] and leads to substantial economic losses [3] due to decreased milk production and reproduction [4] as well as premature culling [3]. While lameness is commonly assessed by trained observers performing visual locomotion scoring of the herd, the procedure is time-consuming and cannot realistically be performed on a regular basis. Hence, dairy farms could benefit from automatic lameness detection.
@@
To date, a number of studies have investigated ways to automate locomotion scoring and lameness detection using camera systems. Video cameras are an attractive sensor for this application as they are relatively inexpensive, non-intrusive, and scale well with large herds. A three-step approach is commonly taken to detect lameness from videos: (1) use computer vision methods to localize body parts of interest, (2) compute one or more locomotion traits from the extracted body parts, and (3) train a classifier to score lameness using the locomotion traits as features. In the past, the body parts were localized using classical computer vision methods such as background subtraction [5, 6, 7, 8]. These methods worked in experimental settings but were sensitive to changes in background and light, making them less applicable in practice. Others placed physical markers (tags or paint marks) on the cows’ body parts and tracked the markers with specialized software [9, 10]. In practical settings, however, physical markers don’t scale well to large herds as they need to be placed on each cow and cleaned regularly to remain visible. More recently, with the emergence of deep neural networks, studies started using deep-learning-based object detection [11, 12, 13, 7] to localize the legs or the back of the cows, object segmentation [14] to extract the body contour from the background, or markerless (i.e., without physical markers) pose estimation [15, 16, 8, 17, 18] to localize multiple body parts in videos. Although they typically require more data than classical approaches, the deep-learning methods cope well with complex background and light conditions and can sometimes even cope with occlusions such as fences [16, 18].
@@
Once localized in the images or video frames, the outline of the spine, for instance, can be used to compute the back posture [6, 19, 20, 21, 22, 13, 7], and the location of the legs to compute the tracking distance [5, 9] or stride length [9, 11, 7]. To the best of our knowledge, almost all studies on lameness detection from videos use only one locomotion trait as a feature to score lameness, and so far, only [23], [17], and [8] combined multiple locomotion traits.
@@
Using the locomotion trait(s) as feature(s), supervised learning classifiers can then be trained to score lameness. In supervised learning, classifiers learn from given examples, also known as ground truth or golden standard. Manual locomotion scores, that is, locomotion scores provided by one or more observers, make up the ground truth of lameness detection classifiers. The subjective nature of manual locomotion scoring is a well-known problem [24] and often leads to low intra- and inter-observer reliability and agreement. However, a classifier can only be as good as its ground truth, so information about the reliability of the locomotion scale is necessary. Yet, observer reliability and agreement are seldom reported, let alone analyzed.
@@
Three critical gaps emerge from the studies discussed so far: (1) the use of obsolete image processing methods remains frequent, (2)few studies combine multiple locomotion traits for lameness classification, and (3) the reliability of the ground truth is seldom reported. This paper addresses the three gaps mentioned above and proposes a non-intrusive and fully automated approach to camera-based lameness detection that includes multiple locomotion traits. Additionally, the code of this paper is open-source [^1] Code available at: https://github.com/hrussel/lameness-detection.
@@
We used videos of walking cows that were scored on a 5-point locomotion scoring scale by four observers. We first reported and discussed the intra- and inter-observer reliability and agreement of the ground truth. We merged scores from the multiple observers to a binary scale. We then trained T-LEAP [16], a deep-learning markerless pose estimation model, to automatically extract the motion of multiple body parts (later referred to as keypoints) from videos of walking cows. The sequences of keypoints were used to compute six locomotion traits that are known to be correlated with locomotion scores [25], namely back posture measurement, head bobbing, tracking distance, stride length, stance duration, and swing duration. Using the locomotion traits mentioned above as input features, we trained multiple machine-learning models to classify the gait as normal or lame. We evaluated the performance of each model and showed the impact of using different combinations of locomotion traits on the lameness classification.
@@ 2.1 Data acquisition
The data were collected in Tilburg, The Netherlands, at a commercial dairy farm whose herd contained about a hundred Holstein-Frisian cows. The data were collected between 9 am and 4 pm on 8 different days between May and July 2019. The cows were filmed from the side while they walked freely through an outdoor passageway. A ZED RGB-D stereo camera [^2] https://www.stereolabs.com/zed-2/ was placed 2 meters above the ground, at 4.5m from the fence of the passageway. The camera directly faced the passageway and recorded in landscape mode at Full-HD (1080p) resolution at 30 frames per second. The recordings were saved into short videos of about 7.6 seconds, which was the average time a cow needed to walk the visible part of the passageway (9.5 meters). The same data acquisition campaign was used by [16] on the same farm. In total, 1101 videos were collected, and a subset of 272 videos were selected according to the following criteria: there was only one cow on the passageway, and the cow walked from the left to the right without distraction or interruption.
@@ 2.2 Locomotion scoring
The locomotion scoring was performed using the 5-point discrete scale described by Sprecher et al. 1997 [26], where a score of 1 corresponds to normal gait, 2 to midly lame, 3 to moderately lame, 4 to lame and 5 to severely lame. Four observers scored the videos: one expert (A) with 20 years of experience in visual locomotion scoring and three observers (B, C, D) with no prior experience in locomotion scoring but with a background in animal science and dairy farming. The inexperienced observers were trained by the expert (A) before the scoring session. During the scoring session, each video was played twice in a row to give enough time to observe the locomotion. To ensure consistency, the observers were asked to give the lowest score if they were hesitating between two scores. All the videos were scored on the same day. After the scoring session, the observers indicated no cow recognition, i.e., that they did not recognize the individual cows that appeared in multiple videos. Table 1 shows the distribution of the scores assigned by the four observers. The distribution of the scores was highly imbalanced and indicated a homogeneous herd, where most cows were distributed throughout the first two levels of the scale (normal, mildly lame), which is typical of herds with a low prevalence of lameness [27].
@@ 2.3 Observers reliability and agreement
Manual locomotion scoring is subjective [28]. Investigating the reliability and agreement between (inter-rater) and among (intra-rater) raters can inform on the quality of the data. Reliability estimates the capability of the raters to differentiate between the different scores, whereas agreement assesses the capability of the raters to assign the same score to the same data point. Reliability was measured with Krippendorff’s $\alpha$ [29] for ordinal values, and agreement was presented as the Percentage of Agreement (PA). The inter-observer and intra-observer measures are reported in Table 2. Note that, as the videos were only scored once, the intra-observer agreement was estimated by comparing the scores between pairs of videos of the same cows recorded at less than 48-hour intervals.
Table 2: Inter-observer (A,B,C,D) and intra-observer reliability and agreement of the locomotion scoring.
Observer (A,B,C,D) A B C D
Krippendorff’s $\alpha$ 0.602 0.611 0.552 0.653 0.585
Percentage of Agreement 55.8 56.4 49.1 60.0 58.2
@@ 2.4 Merging the locomotion scores
In order to create the ground truth for the classification dataset, each sample (video) was assigned a single ground-truth label based on the locomotion scores provided by multiple observers. To achieve this, the scores from the four observers were combined into one value by calculating the rounded-down average of the two observers with the highest intra-observer reliability.
@@
The majority of the studies on lameness detection focus on 2-level (normal, lame) or 3-level (normal, moderately lame, lame) locomotion scales rather than on a 5-level scale [26, 28]. The main motivation for resorting to smaller resolutions in locomotion scales is the distribution of the scores throughout the scale. Severely lame cows are rare, as most get treatment or are culled before they reach this level of lameness [30]. This results in a heavily unbalanced score distribution, most scores being levels 1, 2, and 3. It is then challenging to train a classifier on unbalanced datasets, especially when little examples are available for some classes. As shown in Table 1, the distribution of the locomotion scores was highly imbalanced. Therefore, in order to balance the dataset, we merged the levels of the scale into a binary scale where level 1 indicated a normal gait, and levels 2,3,4 and 5 indicated a lame gait. This resulted in an agreement of 80%, and reliability of 0.590. Note that reliability metrics such as Krippendorff’s $\alpha$ can decrease when the scoring scale is smaller because the chance of agreement is larger.
After combining the scores of the multiple observers and turning to a binary scale, the ground truth for the classification dataset, consisting of 272 videos, contained 143 videos labeled as normal, and 129 videos labeled as lame.
@@ 3 Methods
Our methodology consisted of three main parts: pose estimation, gait features extraction, and lameness classification. These parts are described in detail in the following subsections, and a graphical summary of the methods is provided in Figure 1.
@@ 3.1 Pose estimation
Pose estimation models can be used to predict the position of keypoints (body parts) in images and videos without requiring physical markers. T-LEAP is a recent, deep-learning-based, temporal pose estimation model that was trained to detect keypoints on the body of cows in videos [16]. The model used sequences of successive frames to predict the coordinate of the keypoints, and was shown to perform better than static approaches in the presence of occlusions (such as fences). In this study, we used T-LEAP to extract nine keypoint coordinates from the video frames (Figure 2). In the next paragraphs, we describe the steps necessary for image cropping, pose estimation, and correction.
@@ 3.1.1 Detect-and-crop
The T-LEAP model required the input frames to be square and cropped around the cow’s body. The cows were automatically localized in the video frames using the Faster Region-based Convolutional Neural Network (Faster R-CNN), an object-detection model that returns the coordinates of a bounding box (bbox) around each object of interest (here, cows). We used the Faster R-CNN model (with ResNeXt-101 backbone) trained on the COCO-2017 dataset from the Detectron2 library [31]. The COCO-2017 dataset contained 118K training images with annotations for 80 categories of objects, among which 8014 bounding-box annotations of cows. The Faster R-CNN model from Detectron2 worked out of the box and could detect the cows in our video frames without fine-tuning. Each frame of each video was fed to the object-detection model, which returned a list of bounding boxes, one for each detected cow. For each frame, the bounding box was made square by extending the top and bottom coordinates to match the width while keeping the cow vertically centered. A 100-pixel padding was added to all four sides to ensure that the body of the cow was fully visible in the cropped area. The image was cropped to the coordinates of the extended bounding box and re-scaled to a size of $200\times 200$ pixels. The coordinates of the cropping bounding box were saved to transform the keypoint predictions back to the true coordinates for the video frame.